• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name b. raue

    Log entry time 08:33:50 on October20,2002

    Entry number 88539

    keyword=start of shift

    Start of day shift.
    Began by doing elastics scan to measure spectrometer offsets.
    Finished 9-foil run for setting Ela_H02.
    MCC calls to say they are going down for a while to work on getting us more current.
    Going to be down for 2 hours for machine development.
    Waiting mode...